OpenStudy (hunterk):

2. Michelle made a bead necklace that was 13 inches long. The length of the necklace, in inches, is equal to 3 added to one-third the number of beads, b. How many beads did Michelle use for her necklace?

OpenStudy (herpderp):

13 = 1/3b + 3 10 = 1/3b 30 = b 30 beads